# Efficient Progressive Image Compression with Variance-aware Masking

### v1

@@ Abstract
Learned progressive image compression is gaining momentum as it allows improved image reconstruction as more bits are decoded at the receiver. We propose a progressive image compression method in which an image is first represented as a pair of base-quality and top-quality latent representations. Next, a residual latent representation is encoded as the element-wise difference between the top and base representations. Our scheme enables progressive image compression with element-wise granularity by introducing a masking system that ranks each element of the residual latent representation from most to least important, dividing it into complementary components, which can be transmitted separately to the decoder in order to obtain different reconstruction quality. The masking system does not add further parameters nor complexity. At the receiver, any elements of the top latent representation excluded from the transmitted components can be independently replaced with the mean predicted by the hyperprior architecture, ensuring reliable reconstructions at any intermediate quality level. We also introduced Rate Enhancement Modules (REMs), which refine the estimation of entropy parameters using already decoded components. We obtain results competitive with state-of-the-art competitors, while significantly reducing computational complexity, decoding time, and number of parameters.
@@ 1 Introduction
This article has been accepted for publication at the 2025 IEEE/CVF Winter Conference on Applications of Computer Vision (WACV 2025).
@@
  • We introduced method where, to achieve FGS, complementary portions of a residual latent representation are added to the base one, which represents the lowest level in terms of bitrate.
@@ 4.1 Setup and training details
In the second phase, we refine $g_{s}^{t}$ for 80 epochs with a learning rate of $10^{-4}$, while in the last one we fixed the number of REMs to three to balance performance and complexity, empirically selecting quality checkpoints at ${}{\bar{q}\!=\!\{0.5,7.5,20\}}$ and refining them for 30 epochs. These values, approximately 5%, 40%, and 65% of the Kodak bitstream, ensure that a REM module covers the entire bit range while improving entropy parameter estimation in a sensible portion of the latent space, avoiding excessively high $q$.
@@ 4.3 Model Complexity
We analyze the different ranges defined in Sect. 3.4, to understand how complexity varies when adding additional portions of the bitstream. Thanks to our efficient masking system and highly parallelizable modules, our method is optimal in terms of both time and computational resources versus Jeon et al..
The complexity improvement of our method is substantial, especially for lower bitrates, i.e., where fewer modules, and so fewer iterations, are required. Our gain marginally decreases when increasing the bit rate while always being significantly faster than Jeon et al.. On average, we obtain an improvement of a factor of $2$ both in terms of GFLOPs and decoding time shown in Figures  8. Our method becomes less efficient for higher qualities because of the presence of REMs; in fact, one must first decode the latent representations at the fixed checkpoint quality, which are used as inputs for such modules, and only then can the target quality be achieved. We attribute our superior time performance compared to Jeon et al. to several factors: first, our general architecture, with also the addition of REMs, is smaller than the context model used by Jeon et al. and only needs to be applied once during decoding; furthermore, our masking policy is fast and efficient, whereas Jeon et al. relies on a complex optimization method for RD priority transmission; finally, our algorithm allows one to encode and decode the complementary portions of the final bitstream, making the computation simpler for a specific target quality. Tab. 2 compares the complexity of our method and Jeon et al. on the entire bit range of interest. Our method is better for both GPU and CPU usage, requires fewer parameters, and maintains similar performance at encoding.

### v3

@@ Abstract
Learned progressive image compression is gaining momentum as it allows improved image reconstruction as more bits are decoded at the receiver. We propose a progressive image compression method in which an image is first represented as a pair of base-quality and top-quality latent representations. Next, a residual latent representation is encoded as the element-wise difference between the top and base representations. Our scheme enables progressive image compression with element-wise granularity by introducing a masking system that ranks each element of the residual latent representation from most to least important, dividing it into complementary components, which can be transmitted separately to the decoder in order to obtain different reconstruction quality. The masking system does not add further parameters or complexity. At the receiver, any elements of the top latent representation excluded from the transmitted components can be independently replaced with the mean predicted by the hyperprior architecture, ensuring reliable reconstructions at any intermediate quality level. We also introduced Rate Enhancement Modules (REMs), which refine the estimation of entropy parameters using already decoded components. We obtain results competitive with state-of-the-art competitors, while significantly reducing computational complexity, decoding time, and number of parameters.
@@ 1 Introduction
  • We introduced a method where, to achieve FGS, complementary portions of a residual latent representation are added to the base one, which represents the lowest level in terms of bitrate.
@@ 4.1 Setup and training details
Secondly, we refine $g_{s}^{t}$ for 80 epochs with a learning rate of $10^{-4}$, while in the last one we fixed the number of REMs to three to balance performance and complexity, empirically selecting quality checkpoints at ${}{\bar{q}\!=\!\{0.5,7.5,20\}}$ and refining them for 30 epochs. These values, approximately 5%, 40%, and 65% of the Kodak bitstream, ensure that a REM module covers the entire bit range while improving entropy parameter estimation in a sensible portion of the latent space, avoiding excessively high $q$.
@@ 4.3 Model Complexity
We analyze the different ranges defined in Sect. 3.4, to understand how complexity varies when adding additional portions of the bitstream. With our efficient masking system and parallelizable modules, our method is optimal in time and computational resources compared to Jeon et al..
The complexity improvement of our method is substantial, especially for lower bitrates, i.e., where fewer modules, and so fewer iterations, are required. Our gain marginally decreases when increasing the bit rate while always being significantly faster than Jeon et al.. On average, we obtain an improvement of a factor of $2$ both in terms of GFLOPs and decoding time, as shown in Figures  8.
Our method becomes less efficient at higher qualities due to REMs, as decoding the fixed checkpoint latent representations is necessary before achieving the target quality. We attribute our superior time performance compared to Jeon et al. to several factors: first, our general architecture, with also the addition of REMs, is smaller than the context model used by Jeon et al. and only needs to be applied once during decoding; furthermore, our masking policy is fast and efficient, whereas Jeon et al. relies on a complex optimization method for RD priority transmission; finally, our algorithm allows one to encode and decode the complementary portions of the final bitstream, making the computation simpler for a specific target quality. Tab. 2 compares the complexity of our method and Jeon et al. on the entire bit range of interest. Our method is better for both GPU and CPU usage, requires fewer parameters, and maintains similar performance at encoding.
@@ 5 Conclusions
Acknowledgements
This research was partially funded by Hi!PARIS Center on Data Analytics and Artificial Intelligence.